OpenStudy (anonymous):

Write the equation of the line with a slope of -1/3 and passing through the point (6, -4).

OpenStudy (anonymous):

i have y=-1/3-4

OpenStudy (anonymous):

y = m x + b Plug in the x and y coordinates of the point and the slope m, -1/3. Then solve for b, the y intercept.

OpenStudy (anonymous):

is my answer wrong

OpenStudy (anonymous):

y=-1/3-4 is not the equation of a line. Where is the symbol x in your expression?

OpenStudy (anonymous):

6

OpenStudy (skullpatrol):

Start with $$y=mx+b$$

OpenStudy (skullpatrol):

@excoco have you seen this equation before?

OpenStudy (anonymous):

Another approach is to solve the following expression for y.\[\frac{y-(-4)}{x-6}=-\frac{ 1}{3} \]\[y=-\frac{1}{3}x-2 \]

OpenStudy (anonymous):

@skullpatrol yes i have thats what i started with and got y=-1/3+4

OpenStudy (skullpatrol):

Where is "x"???

OpenStudy (anonymous):

y=-1/3(6)+4

OpenStudy (skullpatrol):



OpenStudy (skullpatrol):